Question 3 Steam expands through a well-insulated turbine from inlet conditions 300°C and 4 MPa with negligible velocity to eSteam expands through a well-insulated turbine from inlet conditions 300c and 4 Mpa with negligible velocity to exit conditions 40m/s and 0.075Mpa. The turbine is operating at steady state and has an exit diameter of 0.6Mpa. Select two different values of exit quality in the range of 0.2 to 0.5 and subsequently plot the power generated by the turbine in kw
Answer #1

Q.3 0-0 ( insulated). Given ws Ti-- Busc P,=&mpa VO } h, = 2961.65 18 h (ov) it Turbine! 1 P2= = a075 MPa hf = 384.36 kJ 1kgTo find mi ma A₂ V2 A, Vi V V2 Now 3 Va= Vf td al Vg - Vf) :: V1.0372X10-3 +0.25 2.217 -1.0372x103) :| = 0.555 milky .. ma Iput all in egn (2 : Wy=17-58 X103 kW Ans.). Plot: Xz us. Wir Wir (kw) 40.902x103 17-58 X103 x2- 0.25 X2=0.45 X2

Please Rate,

Thank you!!!